1.let u=<-2,1,-2>,v=<-3,2,-1> and w=<1,3,5>.compute:

1.1.U*W.

1.2. U*(W*V) and (U*W)*V.


2.find a point- normal form of the equation of the plane passing through p=(1,2,-3) and having n=<2,-1,2> as a normal.



1
Expert's answer
2021-06-25T05:32:34-0400

u = -2i + j - 2k

v = -3i + 2j - k

w = i + 3j + 5k


1.1 u*w = (-2i + j - 2k) * (-3i + 2j - k)


u * w = "\begin{vmatrix}
 i & j &k\\
 -2 & 1& -2\\
 1&3&5\\
\end{vmatrix}" = 11i + 8j -7k



u*w = 11i + 8j -7k ............................equation(1)



1.2 u*(w*v) = ( u.v ) w - ( u.w) v

u*(w*v) = [ (-2i + j - 2k ) . (-3i + 2j - k) ] w - [( -2i + j - 2k).( i + 3j + 5k)] v

u*(w*v) = 10w + 9 v

u*(w*v) = -17i + 38j + 49k






(u*w)*v


from equation(1) we have u*w = 11i + 8j -7k



So, (u*w)*v = ( 11i + 8j -7k ) * (-3i + 2j - k)


(u*w)*v = "\begin{vmatrix}
i&j&k\\
 11 & 8 & -7 \\
 -3 & 2 & -1\\
\end{vmatrix}" = 6i + 32j + 46k



(u*w)*v = 6i + 32j + 46k







2) If a plane passes through point P (x1,y1,z1) and has a normal n = (a, b, c) then the point - normal form of the plane is given by


a(x - x1) + b(y - y1) + c(z - z1) = 0



Here, P = (1,2,-3) and n= 2i - j + 2k


So, the equation of plane in point - normal form is


2(x - 1) - (y - 2) + 2(z +3) = 0


2x - y + 2z + 8 = 0